1. A jeepney driver claims that his average monthly income is Php 3000.00 with a standard deviation

of Php 300.00. A sample of 30 jeepney drivers were surveyed and found that their average monthly 

income is Php 3500.00 with a standard deviation of Php 350.00. Test the hypothesis at 1% level of 

significance


1
Expert's answer
2021-06-27T18:02:44-0400

Solution:

Given, "\mu=3000,n=30,\sigma=350, \bar X=3500"

At 1% significance level and df=29, t=2.462021

"H_0:\mu=3000
\\H_1:\mu\ne3000"

Now, "\dfrac{\bar X-\mu}{\sigma/\sqrt n}=\dfrac{3500-3000}{350/\sqrt {30}}=7.824>2.462021"

"H_0" is rejected.

Thus, "\mu\ne3000"